# Sustainable and Adaptive Growth in Creative Tech

### v1

@@ Abstract.
The creative technology evolves rapidly in both scope and depth, demanding cross-disciplinary expertise and continuous improvement. Although educational programs and other collaborative initiatives enable strong technical and artistic skills, even the most advanced pathways rarely ensure a stable career. Success in these professions often depends on visibility, timing, and self-directed development. As markets shift or technologies change, talents still find themselves displaced. Existing learning paths often fail to connect the skills they teach, leaving learners with fragmented expertise that decays quickly when not continuously applied. The industry demands depth, yet specialization carries risk when tools, pipelines, or roles evolve faster than the expertise built around them. Broad skill sets, by contrast, may increase employability but are easily replaced or rendered obsolete by technological change. CLEAR CORE is a framework for learning and sustaining in creative technology. It integrates two iterative interconnected cycles into a continuous process linking structured education with independent growth as a lifelong, renewable practice that allows professionals to excel amid constant change.
Creative Technology, Skill Development, Career Sustainability, Framework-Based Learning
$c$s: Social and professional topics Computing education programs
@@ 1. Introduction
Highly talented, passionate, and experienced creative technologists are still needed in industries such as gaming and film (TIGA Staff, 2024; Paulo Dias, 2025; Shijia, 2023). Despite the strong industry demand, many professionals shift roles or even move to different industries due to the uncertainties: Some fear unemployment and believe other fields offer more stability, while others, being unemployment, feel they have no choice but to leave the creative tech industry (Sharma and Stol, 2019; Crevoshay et al., 2019; Weststar and Lentini, 2024). When experts in demand are driven away, it raises the question: why does the industry act against its own interests? Many creative tech professionals show strong skills and dedication, evident in platforms like ArtStation (ArtStation Team, 2022) and game jams (GGJ Staff, 2025). With numerous avenues for ongoing collaboration, learning and improvement, their challenges—and those of the industry—seem to lie beyond skill development alone.
@@
Creative technologists should be guided by passion; however, industry realities often dictate choices contrary to it. Both, deep specialization and broad adaptability, is needed, yet pursuing both simultaneously overextend resources, making progress in either direction less efficient (Kelly et al., 2011). Studios look for professionals who excel creatively and technically while understanding how their work fits into the broader pipeline, collaborating effectively, and adapting to new tools and workflows (Sosa et al., 2022). The tension between depth and breadth leaves practitioners in a precarious position: deep specialists risk obsolescence when tools or workflows change (McGuinness et al., 2023), whereas generalists struggle to distinguish themselves in highly competitive markets (Moeran, 2010).
@@
Professionals in the creative technology sector can acquire technical skills effectively; however, translating these abilities into sustained career outcomes remains challenging (Comunian et al., 2011). Investments in learning often exceed the returns (Turegeldinova et al., 2024), and ”there is substantial skill loss with nonpractice or nonuse” (Jr. et al., 1998). Empirical studies suggests that proactive engagement can enhance career outcomes (Chang et al., 2023), and this is widely promoted within creative industry. Yet, when these strategies were widely adopted, competition intensified and returns remained limited, leaving even highly skilled individuals with persistent instability—a reflection of the unequal rewards for professional initiative.
@@
Sustaining professionally in the industry requires visibility (a typical form of proactive engagement), yet research over the past decade indicates that the focus on personal branding often introduced more complications than it supported growth. (Labrecque et al., 2011; Mei et al., 2024). To address these gaps, a structured framework is designed that guides creative technologists in their learning and professional activities, helping them navigating through the known industry challenges. The goal is to support practitioners in building stable, resilient careers while continuing to pursue their passion.
@@ 2. Framework Requirements
To effectively address the challenges faced by creative technologists, the proposed framework must satisfy a set of essential requirements that ensure both professional growth and career sustainability. The following requirements reflect the need to cultivate technical and creative competence, integrate into professional contexts, maintain long-term relevance, and sustain intrinsic motivation:
@@ 3.1.1. Game Jams
  • Skill Evaluation (8): High due to multiple roles and tight deadlines (Borg et al., 2020); slightly reduced because teamwork can buffer individual skill pressure (Saud and Rice, 2024).
  • Credibility (7): Project outcomes are publicly visible and provide networking opportunities but they depend on the ability to effectively collaborate with peers (Pirker et al., 2018).
  • Sustainability (3): Occasional transition to commercial projects exists, but most jams do not generate predictable income; mainly a learning and networking platform (Goddard et al., 2014).
  • Adaptability (6): Opportunities to try new tools exist, but due to time pressure and competition feel most familiar ones are preferred (Lai et al., 2021).
  • Passion (6): Engagement is strong, though social issues often occur (Kultima, 2021).
  • Versatility (7): Rapid prototyping encourages using diverse tools, dependent on personal initiative and team context (Lai et al., 2021).
@@ 3.1.2. Indie Development
  • Skill Evaluation (10): Working solo or in very small teams pushes individuals to their limits; failure has immediate consequences, providing intense learning and skill validation. (Freeman and McNeese, 2019)
  • Credibility (8): Success can earn strong recognition, particularly within niche communities, though visibility is limited among the broader industry (Su, 2023).
  • Sustainability (3): High risk and volatility; income is unpredictable, and maintaining output requires mastery or exceptional luck. Some direct consumer rewards exist, but stability is rare (WifiTalents Team, 2025; Freeman et al., 2023; Su, 2023).
  • Adaptability (7): Indie developers frequently adjust to new tools and workflows, but it is more personal preference rather than formalized evaluation processes (Bruno, 2017).
  • Passion (9): High intrinsic motivation drives projects; autonomy allows personal creative freedom, though external support is limited (Freeman and McNeese, 2019).
  • Versatility (8): Necessity to handle multiple roles encourages broad skill sets, often aligned with industry standards; however these skills may be incomplete or self-taught (Freeman and McNeese, 2019).
@@ 3.1.3. Formal Education
  • Skill Evaluation (7): Structured courses provide strong skill evaluation via assignments and exams (Connor and Marks, 2016), but the assessment often reflects academic standards rather than real-world industry expectations (McGill, 2009). Support from instructors reduces the need to struggle independently, which limits exposure to practical problem-solving (López-Nuñez et al., 2024).
  • Credibility (6): Degrees and school reputation can facilitate industry entry and networking, but they do not guarantee recognition or employment; personal initiative and portfolio quality remain crucial (Jones et al., 2023).
  • Sustainability (7): Formal education, when appropriately focused, can enhance job prospects and earning potential (Czauderna, 2018); however, high costs and delayed entry into the workforce can create financial precarity (Innes, 2025).
  • Adaptability (2): Curricula are still rigid and outdated; learners must follow predefined tools and methods, limiting the development of flexibility and responsiveness to industry changes (Milara and Orduña, 2024).
  • Passion (4): Some motivation comes from collaborative learning and structured guidance (McCord and Exter, 2025), but individual creative choice is constrained by curriculum requirements (Milara and Orduña, 2024).
  • Versatility (6): Education provides a solid foundation and familiarity with industry-standard tools, but exposure to multiple engines or experimental workflows is limited (Jones et al., 2023).
@@ 3.1.4. Employment
  • Skill Evaluation (7): Daily tasks in a studio require consistent skill application, though support from teammates reduces risk of complete failure. Exposure to complex challenges exists, but backup mechanisms moderate the pressure (Saud and Rice, 2024).
  • Credibility (9): In studios, professional recognition develops quickly, and this immediate visibility facilitates networking, making future opportunities more accessible (Espersson et al., 2024). However, outside studio credibility can be dependent on organizational performance and hierarchy.
  • Sustainability (9): Employment provides steady income and benefits. High performers may earn beyond standard rates, but less compared to other industries (National Center for ONET Development, 2025).
  • Adaptability (3): Tool and workflow choices are largely dictated by organizational constraints and licensing, limiting individual flexibility. Some adaptability exists, but it is mostly top-down (Lampel, 2023).
  • Passion (7): Collaborative, skilled teams and tangible end products enhance intrinsic motivation. Personal fulfillment may be limited by team composition, its dynamics or misaligned tasks (Alketbi and Alshurideh, 2023; Weststar and Legault, 2024).
  • Versatility (6): Creative workers can contribute to non-creative industries (Hearn et al., 2014); however, efforts in specialization are often undervalued.
@@ 3.1.5. Portfolio Projects
  • Skill Evaluation (8): Portfolios push practitioners to demonstrate their best work and when aligned with meaningful goals, they can effectively test technical and creative skills but self-directed choices may lead to uneven challenge levels (Richerson et al., 2007).
  • Credibility (7): Portfolios provide visible evidence of ability and can include studio or collaborative work (Watanabe and Misaki, 2025); however, individual contributions may be unclear, and compatibility with team-based work remains uncertain.
  • Sustainability (3): Fully independent portfolio projects carry the risk of commercial non-viability. Even a portfolio career demands multiple income sources to become sustainable (Munnelly, 2022).
@@
  • Passion (8): Practitioners generally pursue projects aligned with personal interest, resulting in motivated and creative output (Keune et al., 2017; Munnelly, 2022); however the excessive external evaluation can undermine the motivation (Amabile, 1979).
  • Versatility (9): Although a portfolio can effectively showcase command of distinct techniques and styles, the individual must choose to display that range (Munnelly, 2022).
@@ 3.3. Observations
In practice, individuals seldom engage with all six methods simultaneously. Each path demands substantial time, focus, and effort, and combining many produces diminishing returns. Humans exhibit limited capacity for effective multitasking (Ophir et al., 2009), which increases the risk of overload when managing multiple creative venues. Therefore, practitioners typically pursue one or two paths—such as employment plus a side project—rather than juggling all methods. This limitation amplifies existing weaknesses: for example, relying primarily on game jams or tutorials yields high adaptability but exposes one to financial and sustainability risks, while focusing solely on employment secures stability but limits creative growth and responsiveness to new technologies.
@@ 6. Conclusion
The CLEAR CORE framework successfully bridges the gap between learning and professional practice for creative technologists. It utilizes dual, interconnected columns for education (CLEAR) and professional application (CORE), leading to a synthesized Outcome layer encompassing Drive, Skill, Credibility, Sustainability, and Mastery that ensures continuous growth and resilience. Initial examples validated the model’s efficacy, demonstrating the closure of the critical gaps in sustainability and adaptability while satisfying all other requirements. The framework’s core achievement is that it ensures all requirements are consistently met—regardless of the practitioner’s chosen method—provided the guiding elements of CLEAR CORE are followed. This guarantees that creative technologists can pursue their passion, excel professionally, and maintain long-term career viability and satisfaction. While this study provides foundational validation, future research will focus on rigorous, quantitative, and comparative longitudinal studies across diverse methodologies to provide deeper, empirical evidence of the framework’s impact in education and professional settings.

### v2

# Sustainable and Adaptive Growth in Computing Education

@@ Abstract.
Computing Education faces significant challenges in equipping graduates with the resilience necessary to remain relevant amid rapid technological change. While existing curricula cultivate computing competencies, they often fail to integrate strategies for sustaining and adapting these skills, leading to reduced career resilience and recurrent industry layoffs. The lack of educational emphasis on sustainability and adaptability amid industry changes perpetuates a vicious cycle: As industries shift, skill fragmentation and decay lead to displacement, which in turn causes further skill degradation. The ongoing deficiency in adaptability and sustainability among learners is reflected in the frequent and intense shifts across the industry. This issue is particularly evident in domains marked by high technological volatility such as computer graphics and game development, where computing concepts— including computational thinking and performance optimization—are uniquely and continuously challenged. To foster sustainable and adaptive growth, this paper introduces, a new framework which addresses the question: How can computing education and professional development be connected to in these volatile sectors? It integrates two iterative, interconnected cycles—an educational and a professional—by linking education with profession to establish a lifelong, renewable practice. This approach allows computing professionals to excel and maintain relevance amid constant changes across their industry.
Creative Technology, Skill Development, Computing Education, Career Sustainability, Framework-Based Learning
$c$s: Social and professional topics Computing education
@@ 1. Introduction
Individuals who combine competencies in computer science with creative practice continue to be highly sought after, especially in creative technology domains—such as gaming and film (tiga2024; combine2025; shijia2023)—which are among the most competitive and volatile professional domains. Despite continued industry growth, many of them shift roles or exit the field altogether due to career instability and uncertainty. Some anticipate long-term unemployment and seek greater stability in other industries, while others, facing layoffs, feel compelled to leave competitive sectors (sharma2019; crevoshay2019; igda2024). This pattern raises a critical question: why do industries that depend on computer science struggle to retain even most capable professionals?
In fields such as computer graphics and game development, professionals consistently demonstrate exceptional technical and creative abilities involving computational thinking, evident through platforms like ArtStation (artstation2022) and game jams (globalgamejam2025). With numerous avenues for ongoing collaboration, learning and improvement, their challenges—and those of the industry—cannot be addressed through technical proficiency alone. Because education’s role extends beyond skill acquisition to preparing learners as industry-ready professionals, it is key to examine which domains face the greatest challenges in transforming students into professionals, and which areas of education are most responsible for enabling that transformation.
Creative-technology, a subdomain of both computer science and art refers to fields that combine computing with creative expression such as game development, computer graphics, interactive media and digital art. The fields in this domain place a greater emphasis on computing education compared to other disciplines (e.g. art), because computing technologies evolve more rapidly and their changes have a much bigger impact on industry practices. While artistic creations often remain timeless, computing tools and frameworks can become obsolete within years, rendering previously valuable expertise less relevant.
@@
Creative technologists—a group of passionate professionals depending heavily on computing education—are expected to possess both deep specialization and broad applicability yet pursuing both simultaneously overextend their resources, making progress in either direction less efficient (kelly2011). Given the highly competitive nature of their industry, even a brief period of skill irrelevance can prompt a career shift or result in unemployment. These professionals are expected to excel creatively and technically while understanding how their work fits into the broader pipeline, collaborating effectively, and adapting to new tools and workflows (sosa2022). The tension between depth and breadth leaves practitioners in a precarious position: deep specialists risk obsolescence when tools or workflows change (mcguinness2023), whereas generalists struggle to distinguish themselves in highly competitive markets (moeran2010).
@@
Acquiring computer science skills is increasingly accessible through diverse pathways, including fully independent learning; however, sustaining and adapting these abilities into stable careers remains a persistent challenge especially in creative technology (comunian01122011). Investments in learning often exceed the returns (fcomm2024), and ”there is substantial skill loss with non-practice or non-use” (arthur1998). Empirical studies suggests that proactive engagement can enhance career outcomes (chang2023), and this is widely promoted within creative industry. Yet, when these strategies were widely adopted, competition intensified and returns remained limited, leaving even highly skilled individuals with persistent instability—a reflection of the unequal rewards for professional initiative.
@@
Sustaining professionally in the industry requires visibility (a typical form of proactive engagement), yet research over the past decade indicates that the focus on personal branding often introduced more complications than it supported growth. (intmar2011; mei2024). To address these gaps, a conceptual, structured framework is designed that guides creative technologists in their learning and professional activities, helping them navigating through the known industry challenges. The goal is to support practitioners in building stable, resilient careers while continuing to pursue their continuous education in computing. By effectively supporting creative technologists, who exemplify the challenges inherent in computing education, this approach can extend its benefits to other computing roles, since the current gaps affect the entire computer science domain.
@@ 2. Framework Requirements
To address the previous mentioned challenges, the proposed framework must satisfy a set of essential requirements that ensure both professional growth and career sustainability. The following requirements reflect the need to cultivate technical and creative competence, integrate into professional contexts, maintain long-term relevance, and sustain intrinsic motivation:
@@ 3.1.1. Game Jams
  • Skill Evaluation (8): High due to multiple roles and tight deadlines (8686138); slightly reduced because teamwork can buffer individual skill pressure (saud2024).
  • Credibility (7): Project outcomes are publicly visible and provide networking opportunities but they depend on the ability to effectively collaborate with peers (10.1145/3196697.3196700).
  • Sustainability (3): Occasional transition to commercial projects exists, but most jams do not generate predictable income; mainly a learning and networking platform (10.1145/2677758.2677778).
  • Adaptability (6): Opportunities to try new tools exist, but due to time pressure and competition feel most familiar ones are preferred (10.1145/3472688.3472689).
  • Passion (6): Engagement is strong, though social issues often occur (10.1145/3472688.3472693).
  • Versatility (7): Rapid prototyping encourages using diverse tools, dependent on personal initiative and team context (10.1145/3472688.3472689).
@@ 3.1.2. Indie Development
  • Skill Evaluation (10): Working solo or in very small teams pushes individuals to their limits; failure has immediate consequences, providing intense learning and skill validation. (freeman2019)
  • Credibility (8): Success can earn strong recognition, particularly within niche communities, though visibility is limited among the broader industry (su2023).
  • Sustainability (3): High risk and volatility; income is unpredictable, and maintaining output requires mastery or exceptional luck. Some direct consumer rewards exist, but stability is rare (wifitalents2025; freeman2023; su2023).
  • Adaptability (7): Indie developers frequently adjust to new tools and workflows, but it is more personal preference rather than formalized evaluation processes (jgeekstudies2017).
  • Passion (9): High intrinsic motivation drives projects; autonomy allows personal creative freedom, though external support is limited (freeman2019).
  • Versatility (8): Necessity to handle multiple roles encourages broad skill sets, often aligned with industry standards; however these skills may be incomplete or self-taught (freeman2019).
@@ 3.1.3. Formal Education
  • Skill Evaluation (7): Structured courses provide strong skill evaluation via assignments and exams (connor2016), but the assessment often reflects academic standards rather than real-world industry expectations (mcgill2009). Support from instructors reduces the need to struggle independently, which limits exposure to practical problem-solving (educsci14111181).
  • Credibility (6): Degrees and school reputation can facilitate industry entry and networking, but they do not guarantee recognition or employment; personal initiative and portfolio quality remain crucial (sign2023).
  • Sustainability (7): Formal education, when appropriately focused, can enhance job prospects and earning potential (czauderna2018); however, high costs and delayed entry into the workforce can create financial precarity (pixlvisn2025).
  • Adaptability (2): Curricula are still rigid and outdated; learners must follow predefined tools and methods, limiting the development of flexibility and responsiveness to industry changes (milara2024).
  • Passion (4): Some motivation comes from collaborative learning and structured guidance (10.1145/3725859), but individual creative choice is constrained by curriculum requirements (milara2024).
  • Versatility (6): Education provides a solid foundation and familiarity with industry-standard tools, but exposure to multiple engines or experimental workflows is limited (sign2023).
@@ 3.1.4. Employment
  • Skill Evaluation (7): Daily tasks in a studio require consistent skill application, though support from teammates reduces risk of complete failure. Exposure to complex challenges exists, but backup mechanisms moderate the pressure (saud2024).
  • Credibility (9): In studios, professional recognition develops quickly, and this immediate visibility facilitates networking, making future opportunities more accessible (espersson2024). However, outside studio credibility can be dependent on organizational performance and hierarchy.
  • Sustainability (9): Employment provides steady income and benefits. High performers may earn beyond standard rates, but less compared to other industries (onet2025).
  • Adaptability (3): Tool and workflow choices are largely dictated by organizational constraints and licensing, limiting individual flexibility. Some adaptability exists, but it is mostly top-down (cgcookies2023).
  • Passion (7): Collaborative, skilled teams and tangible end products enhance intrinsic motivation. Personal fulfillment may be limited by team composition, its dynamics or misaligned tasks (alketbi2023; library994754).
  • Versatility (6): Creative workers can contribute to non-creative industries (hearn2014); however, efforts in specialization are often undervalued.
@@ 3.1.5. Portfolio Projects
  • Skill Evaluation (8): Portfolios push practitioners to demonstrate their best work and when aligned with meaningful goals, they can effectively test technical and creative skills but self-directed choices may lead to uneven challenge levels (4417802).
  • Credibility (7): Portfolios provide visible evidence of ability and can include studio or collaborative work (watanabe2025); however, individual contributions may be unclear, and compatibility with team-based work remains uncertain.
  • Sustainability (3): Fully independent portfolio projects carry the risk of commercial non-viability. Even a portfolio career demands multiple income sources to become sustainable (munnelly2022).
@@
  • Passion (8): Practitioners generally pursue projects aligned with personal interest, resulting in motivated and creative output (keune2017my; munnelly2022); however the excessive external evaluation can undermine the motivation (amabile1979).
  • Versatility (9): Although a portfolio can effectively showcase command of distinct techniques and styles, the individual must choose to display that range (munnelly2022).
@@ 3.3. Observations
In practice, individuals seldom engage with all six methods simultaneously. Each path demands substantial time, focus, and effort, and combining many produces diminishing returns. Humans exhibit limited capacity for effective multitasking (ophir2009), which increases the risk of overload when managing multiple creative venues. Therefore, practitioners typically pursue one or two paths—such as employment plus a side project—rather than juggling all methods. This limitation amplifies existing weaknesses: for example, relying primarily on game jams or tutorials yields high adaptability but exposes one to financial and sustainability risks, while focusing solely on employment secures stability but limits creative growth and responsiveness to new technologies.
@@ 6. Conclusion
The CLEAR CORE framework successfully bridges the gap between learning and professional practice for fields depending on computing education. It utilizes dual, interconnected columns for education (CLEAR) and professional application (CORE), leading to a synthesized Outcome layer encompassing Drive, Skill, Credibility, Sustainability, and Mastery that ensures continuous growth and resilience. Initial examples validated the model’s efficacy, demonstrating the closure of the critical gaps in sustainability and adaptability while satisfying all other requirements. The framework’s core achievement is that it ensures all requirements are consistently met—regardless of the practitioner’s chosen method—provided the guiding elements of CLEAR CORE are followed. This ensures that individuals with a computing education background—particularly creative technologists who encounter the greatest challenges—can pursue their passion, achieve professional excellence, and sustain long-term career viability and satisfaction. While this study provides foundational validation, future research will focus on additional rigorous, quantitative, and comparative longitudinal studies across diverse methodologies to provide deeper, empirical evidence of the framework’s impact in education and professional settings.